Let demand be given by Q = 150 - P + 2Y. This is the same for all problems of this type. Let r = 10%. Let Y = 50 across both periods. Let MC = 0. Let reserves = 200. Let there be a backstop technology available at a constant price of $40. Consider the basic two-period model. What is consumption of the resource in the present?

92.86

105.71

107.14

200.00

213.64
Answer #1

solution: Xan Palsu Given that Q= 150-P+2y. also y=50 Q = 150-P+2659) = 150-pt100 Q = 250-P - By using optimal extraction rul**KINDLY GIVE POSITIVE RATING TO MY ANSWER, THANK YOU IN ADVANCE.